how much pure salt must be added to a 5% salt solution to make 100 pounds of a 24% salt solution?
Answer by ewatrrr(24785)


Hi
Let x represent how much pure salt needs to added to make 100lb of 24% solution
Question states***
"making a 100lb 24% solution
1.0x + .05*(100-x)= .24*100
.95x = .19*100
x = 20 lb of pure salt need to be added.
CHECKING our Answer***
1.0*20lb + .05*80lb = 20lb + 4lb = .24*100 = 24 lbs (pure salt)